# Diagnostic Reasoning in Natural Language: Computational Model and Application

### v1

@@ 4.1 Designing the SCM
We derive the SCM for biomedical text assessment based on expert interviews followed by an independent validation step. Following prior work on modeling complex systems with expert-elicited SCMs Rodrigues et al. (2022); Ashdown and Petrou (2023), we derive a best-effort SCM that is agreeable with the mental paper quality model of domain experts while accounting for subjectivity via the SCM’s background variables. We separately interview two experienced biomedical researchers (postdoc and professor) to understand the criteria for assessing a scientific article in this field. We aggregate and structure their responses to form a step-wise reasoning graph (detailed below), which we revise together with the two experts. Finally, we validate the graph with three PhD students from the same field in a formative study where we ask them to assess a biomedical article following the reasoning graph. Appendix A.1.1 provides details.
The resulting workflow comprises $45$ interdependent paper quality assessment steps terminating in a final verdict on the paper quality. Figure 2 illustrates an excerpt of the workflow. Based on the expert interviews, we identify three key activities during the assessment:
@@
  • for extract steps, humans distill certain information from the previously read text, e.g., listing the results implied by a figure.
@@
We associate each step in the graph with one of these activities, a task prompt, a name, detailed instructions, and an example (see Appendix A.1.2).
The ranges of the endogenous variables are subject to design choice. The assessment of a paper is coupled with its text and figures, requiring a multimodal representation of the inputs. Likewise, in the formative study, we identified that while some steps have boolean values (e.g., "Are the experimental results plausible given the results of related work?"), humans tend to provide detailed, unstructured answers to them. Therefore, we opt for a textual representation of endogenous variables while allowing additional boolean answers for infer nodes and permitting images on the read steps.
@@ 4.2 Populating the SCM
Source data.
We use a sample of openly licensed biomedical manuscripts from bioRxiv [^4] https://www.biorxiv.org/ filtered by recency (06/2021 to 06/2023) and expertise keywords and stratified by topic, world region, and publication status (number of revisions and acceptance status at a peer-reviewed venue). The resulting stratified sample of $200$ papers is further manually filtered by an expert annotator to ensure sufficient expertise and is subject to subselection based on annotator preference, resulting in $22$ papers. See Appendix A.2.1 for more details.
Tools and guidelines
The annotation study was performed in two stages, which we further denote junior and senior based on the level of the participants’ academic expertise. We use similar guidelines, setup, and annotation interfaces for both stages. We turn the graphical workflow into a sequential questionnaire, selecting a permutation of steps that follows the paper structure and adheres to the topological ordering of the graph. This ensures that at any step, the prior answers required as input are already available. We tailor the CARE annotation platform Zyska et al. (2023) for guiding annotators through the assessment while reading the paper. Each step consists of the reasoning step’s name, prompt, detailed description, and an example, plus the prior answers or paper text serving as the input. We provide guidelines and a training session before the study (see Appendix A.2.2).
Annotation study
In the junior stage, the annotation setup was validated within a university course, where $36$ postgraduate students in the biomedical field participated voluntarily. After minor changes to the guidelines and annotator training, the senior stage involved $8$ further doctoral and postdoctoral researchers in the biomedical field, as well as $2$ postgraduate students who received dedicated training in biomedical paper assessment. After the study, we post-processed the junior data to filter out low-quality items (see Appendix A.2.2).
Dataset
The resulting dataset consists of a junior subset with $4$ annotated papers and a senior subset covering $18$ papers. At least three annotators have annotated each paper and workflow step; $11$ papers have annotations by five different annotators. The resulting dataset offers a unique opportunity for the empirical study of human decision-making and reasoning grounded in the text. The following analysis serves as a template to inspect any NL-DAR problem, showcasing the versatility of our framework. The basic statistics of the entire dataset are given in Table 1.
@@
Boolean Decision SCM
To support the analysis, we also construct a condensed version of the workflow that consists only of the boolean variables while maintaining causal ancestry relations. Working on boolean variables makes standard causal estimation methods tractable. We approximate structural equations with Gaussian Process classifiers Rasmussen and Williams (2006). Appendix A.3.2 provides details and reports on the goodness of fit.
@@ 4.3 Analysis
As a first step towards assisting humans in NL-DAR, we aim to learn how humans perform the task. With the populated SCM, we can empirically investigate human decision-making during biomedical paper assessment. While the workflow steps are fixed during the annotation study, the answers to individual steps can and do vary. We evaluate the variation in human answers in the workflow, indicating steps that benefit from AI assistance.
Answer variability
@@
To investigate the variability in the boolean decisions, we compute the inter-annotator agreement on the steps in the boolean decision SCM. Krippendorff’s $\alpha$ per step and across papers is $0.42$, indicating moderate agreement. This is in line with agreement levels reported for peer review ratings Bornmann (2011); Dycke et al. (2021) and typical for subjective annotation tasks. To quantify variability in the natural language answers, we follow recent studies in human answer variability Giulianelli et al. (2023) and quantify the noise induced by background factors in terms of the lexical variability (using lemma overlap), the syntactic variability (using part-of-speech bi-gram overlap), and semantic variability (using cosine similarity on SBERT embeddings Reimers and Gurevych (2019)). We compute the metrics between all annotators per step and paper and then take the average. Figure 3 shows the answer variability distribution for all steps. Notably, the relatively low lexical and syntactical similarity suggests that annotators employ their individual style and register and do not use template language, as has been observed for human-generated explanations in other studies Camburu et al. (2020). The semantic similarity, on average, lies at $0.55$, similar to other open-ended text production tasks, such as story generation Giulianelli et al. (2023). Overall, humans diagnose biomedical papers similarly, but they raise different points. The impact of background variables is reasonably low, suggesting that no important factors were left unmodelled in the SCM.
Sources of variation
@@
We hypothesize that the answer variability changes throughout the workflow graph. Figure 4 shows the semantic similarity distribution for the different node types (infer and extract) and inference steps that explicitly invoke the annotators’ background knowledge (infer-knowledge), such as "From your experience, which plot properties are common for the given plots?". Answers to the steps based on background knowledge deviate the most, whereas extract steps show the highest similarity. The summarization-like extraction steps invoke high consistency. As infer-knowledge steps heavily depend on the individual’s knowledge, prior experience is one major background factor leading to variability. Considering variability as a proxy for step difficulty, AI support during paper assessment on background knowledge shows great potential. We provide additional analysis in Appendix A.3.1, including the propagation dynamics of variability through the causal graph.
@@
Figure 3: Average lexical, syntactic, and semantic similarity per step and paper across all samples. Vertical dotted lines are mean values.
Figure 4: Average semantic similarity by node type. Vertical dotted lines are mean values.
Impact of decisions
@@
The previous analysis focuses on the background variables. We now turn to the explicitly modeled variables of the SCM. We simulate $200$ samples using the boolean decision SCM trained on human data. Using average causal effect estimation (ACE) (Section 3.1), we compare the effect of interventions on different steps on the final verdict. For each node $x$ with a path to the final node $y$, we compute $\textit{ACE}(x,y)$. Table 2 lists the ACEs for selected steps. step33 (Alignment of Conclusions and Research Questions) has the highest positive impact on the final verdict. This means if the findings align well with the posed questions, the final verdict turns more positive. However, the clarity and style of writing also play an important role (step48). As the ACE directly measures the strength of causation (not an associational relation), we can conclude that annotators show a positive bias towards well-phrased papers regardless of the scientific content. This is consistent with findings in scholarly peer review Lee et al. (2013).
Counterfactual analysis
@@
Finally, we explore the use of the SCM for counterfactual analysis. How would one need to change the underlying paper to turn the final verdict positive? We consider scenarios where an annotator gives a negative verdict ($25$ instances across $12$ papers) and study the effect of counterfactual interventions. As studying all possible $2^{18}$ interventions is intractable ($18$ steps in the boolean decision SCM), we focus on steps related to figure quality ($5$) which we call $G$. For all combinations of the figure-related interventions $C=[f_{x}=i_{x}:x\in G]$, we determine $P_{M_{C}}(t=\textit{True})$ for the final verdict $t$. We find that in $40\%$ of the scenarios, the annotator would have kept their negative verdict. For the other $60\%$, an intervention on a single step would have turned their overall assessment: the alignment between the results presented in the figures and the discussion of the results in the text (step12). In other words, the revision of figures would have frequently improved the paper. Overall, this illustrates the impact of specific paper traits on the overall assessment.
@@ 5 LLM Assistance
The NL-DAR framework allows us to decompose complex diagnostic reasoning tasks, opening new opportunities for targeted AI assistance. We illustrate this by using our data to address the following research questions: (A) Do LLMs perform biomedical paper assessments similar to humans? (B) Can LLMs act as effective AI assistants for humans?
Setup
We split the dataset into $20$ test papers ($4089$ answers) and reserve $2$ papers ($282$ answers) for prompt development. We expmeriment with LLama 3 (8B, 32k context window) AI@Meta (2024), Mixtral Jiang et al. (2024), GPT-3.5 turbo Brown et al. (2020), and GPT-4o OpenAI (2023). We design two prompt templates for each LLM, one for the extract steps and one for the infer steps, and perform the experiments zero-shot. Appendix A.4.1 lists details on the LLMs and prompts. We consider BERT-F1-Score Zhang et al. (2020) to measure the general similarity of answers, SummaC Laban et al. (2022), and TRUE Honovich et al. (2022) to measure factual alignment, and compute the F1 score for boolean decisions. We compute the automatic text similarity metrics relative to all annotators’ answers and report the averages.
LLM Diagnosis Performance
We use the SCM created in Section 4 as a scaffolding to run each LLM as a program Schlag et al. (2023); Dohan et al. (2022). We call the LLM per step and provide the answers to the parent steps as input, thereby imposing the SCM’s reasoning structure. As input to read steps, we provide the paper text and replace figures and tables with human descriptions for language-only LLMs, and provide the figures and tables as images to GPT-4o. Table 3 (top) outlines the LLM program performance. The human baseline is estimated using leave-one-out evaluation, which compares the answer of each annotator to the others and averages across all comparisons. We compute the F1 score relative to the majority label for each step as the ground truth.
The GPT models and Mixtral achieve high scores on factual consistency (SummaC and TRUE), marginally above the human baseline. However, BERT-Score ranks LLMs consistently worse. The decision-making performance on the boolean nodes lies below the average human performance by a large margin. We manually examine the discrepancy between high factuality ratings and low decision alignment in Appendix A.4.2, showing that humans reserve negative decisions for extreme cases and focus their answers on a minimal set of facts relevant to the individual assessment. LLMs, on the other hand, tend to restate inputs, split up facts into small components, and make harsher decisions. These factors might affect the factuality metrics favoring the seemingly more extensive and specific LLM answers. Overall, LLMs produce answers with high factual overlap to the average annotator but fall short in decision-making.
Simulating Human-LLM Collaboration
The NL-DAR framework allows exploring simulated human-AI collaboration scenarios. We now assume the human oversees and corrects an LLM at every step, such that the input to the LLM is the human answers from the parent steps or the human-selected paper excerpts for extract steps. We exclude steps based on figures and tables to compare all LLMs. This tests the ability of LLMs to adjust to the personal diagnosis approach of an individual human. Table 3 (bottom, ${io}$-condition) shows the results. As there is a unique answer for each input and step, we compute the F1 score relative to the respective annotator’s decision and not the majority vote as before. The human baseline is a lower-bound estimate for human performance in this scenario. [^5] For a realistic estimate, humans would need to re-annotate steps based on the input of their fellow annotators. For all LLMs, we see a notable performance increase across the board. LLMs effectively adapt to the specific perspective of an annotator as expressed by their personal inputs to each step, generating more aligned answers.
Error Propagation
The stark performance difference between the use of LLMs as a program and human-LLM joint performance suggests error propagation as a factor. To test this, we discard the graph structure entirely and call GPT-4o for each step in isolation using the prompt and the full paper as input. Table 3 shows the results on the last line. We see that the performance drastically increases compared to GPT-4o running as a program but lags behind its performance with human oversight. This strongly suggests that errors on earlier steps in the workflow deprive later steps of information or provide incorrect information for prediction. Overall, our findings show promise for LLMs as diagnostic assistants during paper assessment. More research on their integration with the SCM of an NL-DAR problem is needed. Graph-of-thought Besta et al. (2024a) and similar thought architectures Yao et al. (2023); Besta et al. (2024b) are promising to mitigate the effect of error propagation by backtracking.
BERT-F1 SummaC TRUE F1
human $\textbf{0.799}_{\pm.06}$ $-0.151_{\pm.30}$ $0.151_{\pm.27}$ 0.859
Llama3 Prg $0.752_{\pm.10}$ $-0.274_{\pm.36}$ $0.098_{\pm.30}$ $0.151$
Mixtral Prg $0.761_{\pm.09}$$\textbf{-0.149}_{\pm.26}-$ $0.120_{\pm.32}$ $0.553$
GPT3.5t Prg $0.750_{\pm.11}$ $-0.165_{\pm.33}$ $\textbf{0.159}_{\pm.37}$ $0.513$
GPT4o Prg $0.779_{\pm.07}$ $-0.196_{\pm.29}$ $0.133_{\pm.34}$ $0.696$
@@
GPT4o§ $0.775_{\pm.07}$ $-0.156_{\pm.28}$ $0.187_{\pm.39}$ $0.842$
Table 3: Performance under varying conditions. $\cdot^{\dagger}=$ LLMs run as a program. $\cdot^{io}=$ LLMs receive the human input for each step. GPT4o§ receives prompts and the full paper text without the graph structure. We average over eligible steps and papers, with the standard deviation in subscript. For all metrics, the higher, the better.

### v2

@@ 4.1 Designing the SCM
We derive the workflow for biomedical text assessment based on expert interviews, followed by an independent validation step, opting against the automatic causal discovery of the workflow from data such as peer review reports. Hereby, we ensure close alignment to the true human reasoning process and ensure the quality of the workflow since expert-elicited SCMs are commonly held as the gold standard Zanga et al. (2022); Zhang et al. (2023). Following prior work on modeling complex systems with expert-elicited SCMs Rodrigues et al. (2022); Ashdown and Petrou (2023), we derive a best-effort SCM that is agreeable with the mental paper quality model of domain experts while accounting for subjectivity via the SCM’s background variables. We separately interview two experienced biomedical researchers (postdoc and professor) to understand the criteria for assessing a scientific article in this field. We aggregate and structure their responses to form a step-wise reasoning graph (detailed below), which we revise together with the two experts. Finally, we validate the graph with three PhD students from the same field in a formative study where we ask them to assess a biomedical article using the reasoning graph. Appendix A.3.1 provides details.
The resulting workflow has $45$ interdependent paper quality assessment steps terminating in a final verdict on the paper quality. Fig.2 shows an excerpt of the workflow. Based on the expert interviews, there are three key activities during assessment:
@@
  • for extract steps, humans distill information from the text, e.g., listing the results in a figure.
@@
We assign each workflow step one of these activities, a task prompt, a name, detailed instructions, and an example (see Appendix A.3.2). The variable ranges are up to design choice. Paper assessment involves both text and figures, necessitating a multimodal representation of inputs. In our formative study, we observed that while some steps have boolean values (e.g., "Are the experimental results plausible given the related work?"), humans often provide detailed, unstructured answers. Consequently, we use textual representations for variables, supplementing them with boolean responses for infer nodes and image inputs for read steps.
@@ 4.2 Populating the SCM
Source Data We use openly licensed biomedical manuscripts from bioRxiv [^4] https://www.biorxiv.org/ filtered by recency (06/2021 to 06/2023) and expertise keywords and stratified by topic, world region, and publication status (number of revisions and acceptance status at a peer-reviewed venue). The resulting sample of $200$ papers is filtered manually by an expert annotator to ensure high expertise and is subject to subselection by annotator preference, resulting in $22$ papers. See Appendix A.4.1 for more details.
Tools and Guidelines The annotation study was conducted in two stages, junior and senior, based on participants’ academic expertise. Both stages used the same guidelines, setups, and annotation interfaces. We converted the graphical workflow into a sequential questionnaire, ordered topologically and matching the paper structure. This ensured that prior answers required as inputs were available at each step. The CARE annotation platform Zyska et al. (2023) was customized to guide annotators through the assessment. Each step included the reasoning prompt, an example, and relevant prior answers or paper text as input. Participants received guidelines and a training session before the study (see Appendix A.4.2).
Annotation Study In the junior stage, the annotation setup was validated within a university course, where $36$ postgraduate students in the biomedical field participated voluntarily. After minor changes to the guidelines and annotator training, the senior stage involved $8$ further doctoral and postdoctoral researchers in the biomedical field, as well as $2$ postgraduate students who received dedicated training in biomedical paper assessment. [^5] We report the annotator compensation in Section Ethics. After the study, we post-processed the junior data to filter out low-quality items (see Appendix A.4.2).
Dataset The resulting dataset consists of a junior subset with $4$ annotated papers and a senior subset covering $18$ papers. At least three annotators have annotated each paper and workflow step; $11$ papers have annotations by five different annotators. The amount of annotations per paper aligns with prior work on measuring human variability in a task Giulianelli et al. (2023), which we rely on during analysis in Section 4.3. Additionally, we validate that the dataset size is sufficient for the purposes of our study using statistical analysis and report the details in Appendix A.4.1. The resulting dataset uniquely supports the empirical study of human decision-making and reasoning grounded in text. Our analysis provides a template for examining any STRICTA problem, demonstrating the framework’s versatility. Basic statistics for the dataset are presented in Table 1.
@@
Boolean Decision SCM To support the analysis, we construct a condensed version of the workflow that consists only of the boolean variables while maintaining causal ancestry relations. Working on Boolean variables makes standard causal estimation tractable. We approximate structural equations with Gaussian Process classifiers Rasmussen and Williams (2006). Appendix A.6.2 provides details.
@@ 4.3 Analysis
To assist humans in STRICTA, we first study how they perform the task. Using the populated SCM, we analyze expert decision-making in biomedical paper assessment empirically. While the workflow steps are fixed, answers vary across annotators. We evaluate the human answer variation in the workflow, indicating steps benefiting from AI assistance.
@@
Answer Variability To investigate the variability in the boolean decisions, we compute the inter-annotator agreement on the steps in the boolean decision SCM. Krippendorff’s $\alpha$ per step and across papers is $0.42$, indicating moderate agreement. This is in line with agreement levels reported for peer review ratings Bornmann (2011); Dycke et al. (2021) and typical for subjective annotation tasks. To quantify variability in the natural language answers, we follow Giulianelli et al. (2023) and measure the noise induced by background factors in terms of the lexical variability (using lemma overlap), the syntactic variability (using part-of-speech bi-gram overlap), and semantic variability (using cosine similarity on SBERT embeddings Reimers and Gurevych (2019)). We compute the metrics between all annotators per step and paper, and then take the average. Figure 3 (top) shows the answer variability distribution for all steps. Notably, the low lexical and syntactical similarity suggests that annotators employ their individual style and register and do not use template language, as has been previously observed for human-generated explanations Camburu et al. (2020). The semantic similarity, on average, lies at $0.55$, similar to open-ended text production tasks, such as story generation Giulianelli et al. (2023). Overall, there is a significant overlap between human responses, but they do raise different points. The impact of background variables is reasonably low, suggesting that no important factors were left unmodelled in the SCM.
@@
Sources of Variation We hypothesize that answer variability is not uniform across the workflow. Figure 3 (bottom) shows the semantic similarity distribution for the different node types (infer and extract) and inference steps that explicitly invoke the annotators’ background knowledge (infer-knowledge), such as "From your experience, which plot properties are common for the given plots?". Answers to the steps based on background knowledge deviate the most, whereas extract steps are most similar. The summarization-like extraction steps invoke high consistency. As infer-knowledge steps heavily depend on the individual’s knowledge, prior experience is one major background factor leading to variability. Considering variability as a proxy for step difficulty, AI support during paper assessment on background knowledge shows great potential. Appendix A.6.1 provides additional analysis, including the variability propagation in the graph.
@@
Figure 3: Average lexical, syntactic, and semantic similarity per step and paper (top) and per node type (bottom). Vertical lines are mean values.
@@
Impact of Decisions The previous analysis focuses on the background variables. We now turn to the modeled variables of the SCM. We simulate $200$ samples using the Boolean decision SCM trained on human data. Using average causal effect estimation (ACE) (Section 3.1), we compare the effect of interventions on different steps on the final verdict. For each node $x$ with a path to the final node $y$, we compute $\textit{ACE}(x,y)$. Table 2 lists the ACEs for selected steps. step33 (Alignment of Conclusions and Research Questions) has the highest positive impact on the final verdict. In other words, if the findings align well with the posed questions, the final verdict turns more positive. However, the clarity and style of writing also play an important role (step48). As the ACE directly measures the strength of causation (not an associational relation), we can conclude that annotators show a positive bias towards well-phrased papers regardless of the scientific content. This is consistent with findings in scholarly peer review Lee et al. (2013). This result also shows that the style and writing of the paper directly influence decision making and not only lead to higher variability due to different interpretations of potentially unclear or ambiguous paper contents. This confirms our design decision to include the style of writing as an explicit confounding factor in the workflow.
@@
Counterfactual Analysis Finally, we explore the use of the SCM for counterfactual analysis. How would one need to change the underlying paper to turn the final verdict positive? Focusing on cases where annotators gave a negative verdict ($25$ instances on $12$ papers), we examine counterfactual interventions. As studying all possible $2^{18}$ interventions is intractable ($18$ steps in the boolean decision SCM), we focus on steps related to figure quality ($5$), which we call $G$. For all possible interventions on $G$, $C=[f_{x}=i_{x}:x\in G]$, we determine $P_{M_{C}}(t=\textit{True})$ for the final verdict $t$. In $40\%$ of the cases, the annotator would have maintained their negative verdict. In the remaining $60\%$, an intervention on a single step – the alignment between the figures and the text discussion (step12) – would have changed the overall verdict. In other words, the revision of figures would have frequently improved the paper. Overall, this illustrates the impact of specific paper traits on the final verdict.
@@ 5 LLM Assistance
The STRICTA framework decomposes complex document assessment tasks, opening new opportunities for targeted assistance and interpretable AI. We investigate how LLMs can support STRICTA, by testing their abilities on the collected dataset. Following Definition 1, the exact reasoning task depends on the data provided to the LLM. For our experiments, we focus on an abductive setting where the final verdict is given and all intermediate reasoning steps need to be reconstructed based on the paper. [^6] We experiment with multiple reasoning settings in Appendix A.7.2. The results are consistent. This models the case where humans have full agency over the verdict but receive assistance to find a valid explanation for it. Specifically, we evaluate LLM alignment with human answers on intermediate steps and in a simulated assistance scenario. We highlight that our goal here is not to introduce a new NLP method, but to explore the boundaries and opportunities of existing LLMs when applied to STRICTA problems. Nonetheless, our experiments shine a new light on LLMs for causal reasoning, and suggest that a fixed reasoning structure may help address some of their inherent limitations Zečević et al. (2023).
Setup We split the dataset into $20$ test papers ($4089$ answers) and reserve $2$ papers ($282$ answers) for prompt development. We experiment with LLama 3 (8B, 32k context window) AI@Meta (2024), Mixtral Jiang et al. (2024), GPT-3.5 turbo Brown et al. (2020), and GPT-4o OpenAI (2023). Appendix A.7.1 lists details on the LLMs and prompts. All prompts are designed zero-shot. We assess LLM outputs using automatic text similarity metrics to measure alignment with human answers across multiple dimensions. While being imperfect Sottana et al. (2023), these metrics capture individual alignment in a human-AI collaboration setting, consistent with prior work on peer review generation comparing automated and human-written reviews Du et al. (2024). We use BERT-F1 Zhang et al. (2020) for semantic similarity, SummaC Laban et al. (2022) and TRUE Honovich et al. (2022) for factual alignment, and F1 score for boolean decisions. Each annotator’s workflow execution is treated as an individual problem instance, with similarity metrics computed per annotator and averaged across all instances. For validation, we complement automatic metrics with a human evaluation on a subset of the data (Appendix A.7.3). Results show moderate agreement among humans – typical for challenging tasks – and a moderate but significant correlation between metrics and human judgment (between $\rho=0.37$ and $\rho=0.54$ for BERT-F1, and $\rho=-0.22$ and $\rho=0.55$ for SummaC, depending on the step type) at similar levels as related automatic evaluation metrics Fu et al. (2024). Yet, these findings highlight the need for more metric development in the future for a targeted evaluation of individual steps.
LLM Structured Reasoning The target task involves answering each step in the workflow to ensure that the given verdict is validly derived from the given paper text. We define a baseline LLM architecture for this process. First, we use the workflow from Section 4 as a scaffold, running each LLM as a program Schlag et al. (2023); Dohan et al. (2022). The LLM is called step-by-step, with the answers from parent steps provided as inputs, thereby enforcing the workflow’s reasoning structure. However, at this stage, the responses are not conditioned on the verdict being explained. To address this, in the second step, we adopt the self-refinement paradigm Madaan et al. (2024), incorporating previous model responses, the graph structure encoded with incident encoding Fatemi et al. (2024), and the true verdict to refine the answers. We design tailored prompts for each workflow step type (extract, infer) and for the feedback and refinement stages. We terminate the feedback-refinement loop after one iteration. For read steps, the input consists of the paper text, with figures and tables replaced by annotator descriptions for language-only LLMs. For GPT-4o, we provide figures and tables as images. Table 3 (top) summarizes the performance of the LLM program. We estimate the human baseline using leave-one-out evaluation, where each annotator’s answers for a paper are compared to those of their peers, averaged across all comparisons. This approach provides a conservative lower-bound estimate of human performance. [^7] An accurate estimate would require humans to re-annotate steps using inputs from their fellow annotators. We calculate the F1 score based on individual annotator responses. The GPT models and Mixtral achieve high scores on factual consistency (SummaC and TRUE), marginally above the human baseline. However, BERT-Score ranks LLMs consistently lower. The decision-making performance on the Boolean nodes lies below the average human performance by a large margin. We manually examine the discrepancy between high factual and low decision alignment in Appendix A.7.4, showing that humans reserve negative decisions for extreme cases and focus their answers on a minimal set of facts relevant to the individual assessment. However, LLMs tend to restate inputs, split up facts into small components, and make harsher decisions. These factors might affect the factual alignment metrics, favoring the seemingly more extensive and specific LLM answers. This is also supported by the relatively low factual alignment of human answers to extract steps with the original paper text compared to the more extensive LLM responses (details in Appendix A.7.2). While LLMs generate responses with high factual overlap compared to the average annotator, they underperform in decision-making. This indicates that LLMs weigh factors differently than humans during paper assessment, highlighting the risks of using LLMs without human oversight in STRICTA applications.
Simulating Human-LLM Collaboration The STRICTA framework allows exploring simulated human-AI collaboration scenarios. Here, we assume the human oversees and corrects the LLM at each step, providing the input as human-generated answers from parent steps or the paper for extract steps. We exclude steps based on figures and tables to compare all LLMs. This setup evaluates the ability of LLMs to adapt to the individual’s stepwise reasoning. We exclude self-refinement to focus on model performance at individual steps. Table 3 (bottom, ${io}$-condition) shows the results. All LLMs demonstrate a notable performance increase, effectively adapting to the annotator’s perspective based on their step-by-step inputs and producing more aligned responses. We further investigate the reasons for these performance differences, focusing on error propagation, in Appendix A.7.2. Overall, our results show the potential of human-AI collaboration for STRICTA, but further research is needed to integrate this approach with the SCM of a STRICTA problem. Promising directions to mitigate error propagation include Graph-of-Thought Besta et al. (2024a) and similar architectures Yao et al. (2023); Besta et al. (2024b), which enable backtracking during reasoning.
BERT-F1 $\uparrow$ SummaC $\uparrow$ TRUE $\uparrow$ F1 $\uparrow$
human∗† $\textbf{0.799}_{\pm.06}$ $-0.151_{\pm.30}$ $0.151_{\pm.27}$ 0.801
Llama3 Prg $0.752_{\pm.10}$ $-0.274_{\pm.36}$ $0.098_{\pm.30}$ $0.170$
Mixtral Prg $0.761_{\pm.09}$$\textbf{-0.149}_{\pm.26}-$ $0.120_{\pm.32}$ $0.559$
GPT3.5t Prg $0.759_{\pm.10}$ $-0.178_{\pm.35}$ $\textbf{0.163}_{\pm.37}$ $0.531$
GPT4o Prg $0.780_{\pm.07}$ $-0.186_{\pm.30}$ $0.139_{\pm.35}$ $0.720$
@@
GPT4o§ $0.776_{\pm.07}$ $-0.154_{\pm 0.28}$ $0.188_{\pm.39}$ $0.828$
Table 3: Performance under varying conditions. $\cdot^{\dagger}=$ LLMs as a program. $\cdot^{io}=$ with human oversight. GPT4o§ does not use the graph. We average over steps and papers, with the standard deviation in subscript. SummaC lies in $[-1,1]$, the other metrics in $[0,1]$.
6 Transfer and Future Work
We design STRICTA as a general domain-independent framework for critical text assessment. To further validate the application to other tasks besides the case study, we additionally test the generality of the STRICTA framework in a transfer experiment to paper assessment in the natural language processing domain. We adapt the biomedical assessment workflow for the NLP domain with input from two experts (PhD and Postdoc in NLP) and conduct a formative study among PhD students. We provide the detailed results in Appendix A.5. Overall, participants found the workflow appropriate, supporting the notion that the framework can be transferred to other domains with reasonable effort. Since the scope of this transfer study lies in the validation of the adapted workflow, we do not undertake quality assurance measures, making the output dataset not suitable for evaluation or analysis. Additional validation of STRICTA  for other text assessment tasks such as essay grading, fact checking, or Wikipedia article quality assessment is important future work. This entails performing the three steps of the STRICTA framework (Section 3.3): (1) defining a workflow, (2) collecting data, and (3) using the resulting SCM and data for analysis and evaluation. Following up on our assistance experiments, humans and LLMs can be employed to jointly generate data Li et al. (2023) for speed-up. To illustrate the process of transferring STRICTA to other text assessment tasks, we describe how to instantiate STRICTA for automatic essay scoring in Appendix A.2.3.